OpenStudy (weasel123):

anyone know how to put g(x)=2x-1 in a composite function?

OpenStudy (danjs):

sure, what other function or other(s) you wan to use

OpenStudy (weasel123):

f(x)= x+1 2 like this?

OpenStudy (danjs):

f(x) = x + 1 g(x) = 2x + 1

OpenStudy (danjs):

composite functions chain 2 or more functions together. the value of one of the function depends on the value for the other function for any input value.. for example, think of another function w(x), and it is the composite function of f and g w(x) = f [ g(x) ] and f(x) = x + 1 g(x) = 2x + 1 ----------------

OpenStudy (danjs):

individually , f(x) and g(x) are both functions of x directly but, f [ g(x)] is 2 functioins chained together say x=1 you have to evalueate g(1) first, g(1) = 2*1 + 1 = 3 the f is now a function of g f(g(1)) = f(3) = 3+1 = 4

OpenStudy (danjs):

basically , the output of one function goes directly to the input of the next function w(x) = f [ g(x) ] g(x) will define what value f needs to evaluate f(g(x)) = f(2x+1) = (2x + 1) + 1